opdaxy:

question: a 12,000 gallon pool is being filled at a rate of 40 gallons per minute. At this rate how many minutes will it take to fill this pool 3/4th full?

opdaxy:

do you know xiogonz?

XioGonz:

First we have to find what 3/4 of the pool is equal to. So we have to multiply it by 12,000 12000* 3/4 Which equals to =9000 gallons If the rate = 40 gallons per minute Then: 9000/40 =225 minutes Now we can see that it would take 225 minutes to fill up 3/4 of the pool.

opdaxy:

hey

XioGonz:

?

opdaxy:

can u help me on another one?

XioGonz:

Sure simply post it in the Q section and tag me there like @xiogonz

opdaxy:

k

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz First we have to find what 3/4 of the pool is equal to. good job @XioGonz So we have to multiply it by 12,000 12000* 3/4 Which equals to =9000 gallons If the rate = 40 gallons per minute Then: 9000/40 =225 minutes Now we can see that it would take 225 minutes to fill up 3/4 of the pool. \(\color{#0cbb34}{\text{End of Quote}}\)

XioGonz:

Thank you @jhonyy9

jhonyy9:

np yw